### v1

# Big Tech influence over AI research revisited: memetic analysis of attribution of ideas to affiliation

Abstract
There exists a growing discourse around the domination of Big Tech on the landscape of artificial intelligence (AI) research, yet our comprehension of this phenomenon remains cursory. This paper aims to broaden and deepen our understanding of Big Tech’s reach and power within AI research. It highlights the dominance not merely in terms of sheer publication volume but rather in the propagation of new ideas or memes. Current studies often oversimplify the concept of influence to the share of affiliations in academic papers, typically sourced from limited databases such as arXiv or specific academic conferences.
The main goal of this paper is to unravel the specific nuances of such influence, determining which AI ideas are predominantly driven by Big Tech entities. By employing network and memetic analysis on AI-oriented paper abstracts and their citation network, we are able to grasp a deeper insight into this phenomenon. By utilizing two databases: OpenAlex and S2ORC, we are able to perform such analysis on a much bigger scale than previous attempts.
Our findings suggest, that while Big Tech-affiliated papers are disproportionately more cited in some areas, the most cited papers are those affiliated with both Big Tech and Academia. Focusing on the most contagious memes, their attribution to specific affiliation groups (Big Tech, Academia, mixed affiliation) seems to be equally distributed between those three groups. This suggests that the notion of Big Tech domination over AI research is oversimplified in the discourse.
Ultimately, this more nuanced understanding of Big Tech’s and Academia’s influence could inform a more symbiotic alliance between these stakeholders which would better serve the dual goals of societal welfare and the scientific integrity of AI research.
keywords:
[…]
1 Introduction
Artificial Intelligence (AI) research is often seen as dominated by Big Tech companies. These companies, with their immense computing resources and access to vast amounts of data, have undoubtedly influenced the development of the field. However, while this heavy influence has been beneficial in certain aspects, it has also raised concerns that the strong position of Big Tech could influence the direction and character of AI, resulting in significant losses for society and science. One of the issues raised is the reproducibility of Big Tech’s research (Ebell et al., 2021). Given their proprietary datasets and custom software, it is often difficult to independently replicate their studies. Conflict of interest is another issue that has been raised by researchers (Abdalla and Abdalla, 2021; Young et al., 2022; Hagendorff and Meding, 2021). One notable case that exposed these concerns was that of Dr. Timnit Gebru, formerly part of Google’s AI ethics team. Her dismissal highlighted the conflicts of interest that can arise when industry-aligned research agendas collide with ethical considerations. There is also evidence that the longer researchers remain in the industry, the more their high-impact work tends to become privatized, limiting the dissemination of knowledge that could otherwise benefit the academic community (Jurowetzki et al., 2021).
While there are many concerns about potential threats to research integrity due to this influence, it is equally important to recognize the positive impact of Big Tech companies. First and foremost, their computational resources and data, previously inaccessible on such a scale, provide unprecedented opportunities for the discovery and development of novel Machine Learning (ML) algorithms. In addition, Big Tech–Academia collaboration may benefit scientific endeavors for reasons other than access to resources. Evans (2010) argues that the industry’s relative lack of interest, in theory, could result in their academic partners being more likely to produce novel and theoretically unexpected experiments than academics working by themselves.
The matters mentioned above indicate that the influence of Big Tech (and more broadly the private sector, also referred to as company in our paper) on AI research needs to be carefully examined. Several such studies in this regard exist, including analysis of Big Tech’s funding of academic AI researchers (Abdalla and Abdalla, 2021), changes in the proportion of Big Tech–affiliated papers at top conferences (Ahmed and Wahed, 2020), the flow of researchers from Academia to the private sector (Jurowetzki et al., 2021), the focus of the private sector on specific sub-fields of AI (Klinger et al., 2020), and the comparison of values encoded in the papers between the private sector and Academia (Birhane et al., 2021).
Unfortunately, approaches to this topic mentioned above are limited. The methods used so far to quantify Big Tech’s dominance tend to oversimplify the concept of influence by perceiving it as the share of affiliations in papers regarding a specific topic. Methods are limited to tools such as topic modeling and keyword analysis. Current research also suffers from a lack of diversity in data sources, relying heavily on preprint servers such as arXiv or specific conferences to collect papers. In the previous studies mentioned above, the association of papers with affiliations was binary. That is, each paper was associated with either Big Tech (or the private sector in general) or Academia, depending on the first author or the proportion of affiliations in the paper. This limits conclusions that could be drawn from studies, as the Company–Academia collaboration in the area of AI is widespread.
All of the above limitations, combined with strong statements (e.g. researchers on this topic do not shy away from comparing Big Tech to Big Tobacco, Abdalla and Abdalla, 2021), could result in a limited, and possibly biased view of the matter, which could increase the tension between Big Tech and Academia, resulting in losses for science and society. Through our study, we aim to facilitate a more nuanced understanding of the dynamics between Big Tech and Academia and to highlight potential strengths and weaknesses within this relationship. We critically evaluate existing methodologies, advocate for more holistic measures of influence, and explore the implications of this dominance for the field of AI. We hope that this work will foster further discourse about the symbiosis between Academia and industry, and help to shape a more balanced AI research landscape.
In this study, along with a standard network analysis aimed at quantifying the influence of Big Tech and Academia on the AI papers citation network, we performed a memetic analysis. A meme is understood here as a piece of an idea that is transmitted through culture. In the context of research papers, this transmission occurs through citation. Using meme score, which is the quantification of the replicative power of a specific meme, we measure the spread of specific ideas in AI research. In addition, we quantify the probability of specific memes being replicated, conditioned on the affiliation of the authors of the paper containing the meme. This approach allows us to understand how the "spreading power" of specific ideas depends on affiliation groups (Big Tech, Academia, Companies, and mixed), which sheds light on which ideas Big Tech has more influence on spreading. Moreover, we investigated how papers with joint Big Tech–Academia affiliations differ from papers authored purely by authors affiliated with one category. In summary, we address the following research questions:
1.
Do papers affiliated with Companies and papers affiliated with both Big Tech and Academia differ in citation distributions from papers affiliated only with Big Tech or Academia?
2.
What ideas in AI research are the most contagious?
3.
Does the contagiousness of a meme differ depending on the affiliation of paper authors?
4.
What ideas are more contagious when discussed by Big Tech?
2 Related Work
2.1 Big tech influence over AI research
Several studies have attempted to quantify Big Tech’s influence in AI research, using a variety of methods and data sources. Abdalla and Abdalla (2021) examine the funding provided by Big Tech to academic AI researchers. The study identifies a recurring pattern where private companies increase their support for academic institutions when their public image declines, often due to media incidents such as the Cambridge Analytica scandal. The authors compare this funding pattern to the tactics employed by the tobacco industry. The study reveals that 59% of papers published in top journals that address the ethical and societal implications of AI include at least one author with financial ties to a Big Tech company. Ahmed and Wahed (2020) examine the changing dynamics of participation in major AI conferences following the rise of deep learning in 2012. The authors analyze 171,394 papers from 57 computer science conferences. They find an increase in the participation of large technology companies, particularly since 2012. In addition, the paper uses term frequency analysis of abstracts to uncover distinct research areas among different organizations.
Jurowetzki et al. (2021) use bibliographic data to measure the flow of researchers from Academia to industry and examine the factors driving it. They find that 25% of the AI researchers moved to industry from institutions at the top 5 of the Nature Index. This observation suggests that industry tends to attract AI researchers from elite institutions, possibly reflecting a search for current and potential superstar talent or a narrow focus on high-prestige sources of talent.
Zhang et al. (2019) employs altmetrics to assess the influence of AI publications, revealing an increased public interest in AI research findings since 2011. However, the literature suggests that altmetrics may not be suited to measuring societal impact Bornmann (2014).
The recent article of Färber and Tampakis (2023) strikes to measure the relationship between authors’ affiliations with the private sector and the article’s popularity, measured by citations and attention score. They perform the keyword analysis with respect to the extent of association with the private sector. Their quantitative analysis shows the domination of the private sector in the AI research domain.
Krieger et al. (2021) note an upward trend in private sector publications in fundamental research journals as opposed to those centered on applied research. They also highlight a rise in collaborative publications with academic entities in contrast to solo publishing efforts.
Several studies have tried to analyze the content of AI papers to study the influence of Big Tech. These include topic modeling of the abstracts. Klinger et al. (2018) combine data from arXiv, GRID (Global Research Identifier) and MAG (Microsoft Academic Graph) to create a geocoded dataset of research activity in computer science disciplines. They identify deep learning papers using topic modeling. Finally, the authors measure the relatedness of computer science subjects based on their co-occurrence in arXiv papers. Klinger et al. (2020) analyze the field of study assigned to the papers based on arXiv. They find that companies focus more on applications of deep learning and on research advancing the computational infrastructure. AI techniques outside deep learning and broader AI applications, are of less interest to the private sector. Birhane et al. (2021) perform textual analysis in order to extract values encoded in papers. They analyze affiliations and funding sources in papers and find that the presence of Big Tech is increasing.
However, there is a significant gap in our understanding of Big Tech’s influence. All of the above studies have at least one of the following limitations.
Firstly, none of the studies described above link the content of papers to citations in any way. Therefore, the resulting purely fraction-based analysis of content does not measure the spread of the ideas, but merely their prevalence. Quantifying both the spread and prevalence of specific ideas by using meme score allows us to find the most contagious ideas, called memes. The impact of affiliation groups on the contagiousness of each meme could be then modeled, resulting in a more fine-grained view of the influence of Big Tech, Academia, and other groups. Moreover, this approach allows for a comparison of the ideas on which contagiousness is most affected by Big Tech and Academia, which could provide insight into areas where Big Tech or Academia has the most influence.
Secondly, they simplify the notion of influence to the share of affiliations in papers regarding a specific topic. They use the share of papers affiliated with a particular group only (with the exception of Färber and Tampakis (2023)) that get published at prestigious conferences and in top journals as an indicator of the prevalence of Big Tech in AI research and interpret it as a proxy for influence measure. However, this approach does not take into account the number of the paper’s citations, which is a more reliable proxy for the popularity of the paper and thus closer to measuring actual influence on the research field.
The third limitation is the quantity and diversity of data. The use of manual annotation limits the scope of the captured categories. Similarly, using only arXiv or specific conferences as the source of the papers limits the representativeness of the data.
2.1.1 Operationalization of the concept of Big Tech papers
Author affiliation is a key characteristic associated with the author rather than the research paper itself, although even a simple quantitative approach (i.e., the number of authors) has been proven to be connected to the impact exerted by an article (Sienkiewicz and Altmann, 2016). When examining the impact of Big Tech companies on scientific research, it becomes necessary to determine how to classify papers that are not exclusively authored by individuals affiliated with Academia or Big Tech. Existing approaches (Klinger et al., 2020; Ahmed and Wahed, 2020; Birhane et al., 2021) typically assume that any paper with at least one author affiliated with a Big Tech company should be categorized as a Big Tech paper. Alternatively, they consider the affiliation of the first author as the defining criterion.
However, these approaches oversimplify a more complex situation. Grouping papers authored exclusively by researchers from the private sector, along with papers primarily written by academics, one of whom may have a dual affiliation, seems counter-intuitive. It has been observed that researchers associated with Big Tech companies benefit from access to computational resources, tools, and datasets that would otherwise be unavailable to them (Whittaker, 2021; Jurowetzki et al., 2021). Considering only the first author’s affiliation would lead to overlooking these phenomena.
Another approach is to look separately at papers that result from collaborations between Big Tech companies and other institutions. Articles describing such collaborations tend to have different characteristics in terms of format and tone compared to the research papers mentioned above, and focus on the positive aspects of such collaborations (Popkin, 2019). In recent work, (Färber and Tampakis, 2023) separatly analyzed papers co-authored by Academia and company-affiliated researchers. To the best of our knowledge, no analyses have specifically examined the influence and topics of interest reflected in papers resulting from collaborations between Big Tech and academia.
Furthermore, we are not aware of any analysis that compares Big Tech-affiliated papers to those affiliated with the private sector in general. This limits the ability to assess whether found characteristics are specific to Big Tech, or are phenomena visible in all research originating from the private sector. In our study, we examine the differences between several affiliation groups: purely Big Tech, purely academic, private sector as a whole, and papers affiliated with both Big Tech/private sector and Academia.
2.2 Memetic analysis
Several measures can characterize the spread of ideas in science (Wagner et al., 2011; Xu et al., 2018) or in business researcg (Wu et al., 2017). A straightforward method to identify ideas within a set of documents is to examine their frequency. Techniques like tf-idf (term frequency–inverse document frequency) allow the discovery of potential keywords. Topic modeling methods, such as Latent Dirichlet Allocation (LDA) or BERTopic (Grootendorst, 2022), help to understand the distribution of ideas consisting of multiple phrases across a document set. Another option is to focus on topical clusters, which can be built from bibliographic coupling and co-citation networks (Liu et al., 2017, 2019). One can use a citations cascade (Min et al., 2021) or chains of citations (della Briotta Parolo et al., 2020), which reveal a higher-order (i.e., reaching further than just the neighborhood of the node) temporal relationship between the papers.
One perspective on the ideas’ propagation is looking at it through the lens of the evolutionary theory of science (Kantorovich, 2014). Central to this theory is the concept of meme, first introduced in a broader context by Dawkins (1976). A meme is considered a cultural analog to a biological gene, a unit of information that replicates itself, mutates, and undergoes selection in the evolutionary process. Analogous to genes in biology, memes are pieces of information transmitted between individuals within a given culture.
Kuhn et al. (2014) proposed an operationalization of the concept of a meme in the context of scientific knowledge by introducing the notion of a meme score. The meme score measures how much a given term is a meme by analyzing the citation network. It takes into account the term’s frequency, the probability of its transmission through citation, and the probability of it emerging independently. In this operationalization of memes, a term exhibits more memetic quality, the higher the meme score of that term is. Since its proposal, the meme score has been used in several studies. It has been used to investigate how the gender of researchers influences their positions in the scientific community (Araújo and Fontainha, 2018), to identify scientific and technological trajectories of ideas in paper and patent networks (Sun and Ding, 2018), and to explore diffusion cascades of ideas (Mao et al., 2020).
3 Materials and methods
3.1 Meme score
The meme score was introduced by Kuhn et al. (2014) as a measure of the ’contagiousness’ of meme.
The meme score for the given meme is calculated by multiplying the relative frequency of a term by its sticking factor and then dividing it by its sparking factor. The quotient of the sticking factor and the sparking factor is labeled as a propagation score. Both measures are briefly described below.
The sticking factor can be interpreted as the probability of the meme being transmitted from one paper to the other paper. It corresponds to the probability that the meme appears in the abstract of the given paper, given that it appears in at least one of the abstracts of the papers cited by it. It is calculated by dividing the number of papers that contain the meme and cite at least one paper with this meme by the number of papers that cite at least one paper with this meme: by .
The sparking factor can be interpreted as the probability of the meme appearing without being present in any of the cited papers - how likely it is that the meme will be mentioned in the abstract, given that it does not appear in any of the cited papers’ abstracts. It is calculated by dividing the number of papers with the given meme that do not cite any papers with this meme by - all papers that do not cite any papers with the given meme.
The formula for the meme score can be written as follows:
where is the number of papers with the meme , is the number of all memes in the dataset. is the number of papers that contain the meme and cite at least one paper with the meme and is the number of papers that cite at least one paper with the meme. is the number of papers containing the meme, but not citing a paper with the meme and is the number of papers that do not cite a paper with the meme.
From the three components of the meme score it is mainly the sticking factor that measures meme’s contagiousness. The other factors control if the phrase is not merely used as a part of a natural language or if a phrase is popular enough. Due to this, the sticking factor on its own could serve as a weaker measure of contagiousness.
3.2 Conditioned sticking factor
To measure how probable a meme is to spread under a given affiliation, we introduce a conditioned sticking factor, which is a modified version of the meme score introduced by Kuhn. A similar approach was proposed in the paper of Araújo and Fontainha (2018), where the authors attempted to condition the meme score on the main author’s gender. In contrast, in our approach, we condition only the sticking factor on the authors’ affiliations. The conditioned sticking factor is calculated as
The conditioned sticking factor divides the number of papers that have meme and cite at least one paper with this meme and given affiliation by the number of papers that cite at least one paper with given meme and affiliation . It can be interpreted as the probability of the meme transitioning during citation under the condition that the meme appeared in the paper with a given affiliation.
In contrast to the earlier approach of Araújo and Fontainha (2018), we have decided not to condition the sparking factor. The original purpose of the sparking factor is to control how often the word appears without any influence and limit the appearance of the most popular words that do not convey specific ideas. The sparking factor, conditioned analogously to the sticking factor, would correspond to the probability that the meme appears in a paper without being present in any cited papers with a given affiliation. However, the meme can appear in the other cited papers with different affiliations. In this way, the conditioned sparking factor would lose its original purpose of excluding phrases not conveying ideas without providing us with any specific additional information. For this reason, we decided to condition only the sticking factor. Finally, once , i.e., the meme is relatively common but rarely transmitted through the citation network, the meme score can be approximated by the sticking factor. Due to this, we focus primarily on the relationships between conditioned sticking factors for different groups. Figure 1 presents a toy example to explain how the sticking factor, sparking factor, meme score and conditioned sticking factor are calculated.
3.3 Papers dataset and processing
We gathered 557 681 articles from the S2ORC database, by searching for phrases proposed by Liu et al. (2021) (for details see A) in the titles and abstracts of the papers (all in English). We then performed exploratory data analysis. We found out that 138 878 records had no DOI identifiers, which we used to match the papers against the OpenAlex database. These papers were removed from our dataset. Next, we discovered that 146 525 records had no information about citations – such papers like this were useless for citation network analysis, so we decided to remove them from our dataset. Finally, we ended up with 166 455 records that had affiliations matched with OpenAlex.
OpenAlex provides information about the type of institution – it categorizes them into the following categories: company, education, government, facility, healthcare, nonprofit, and other (Table 2). By analyzing these categories, we decided to merge education, facility, and government into Academia. We also analyzed the company category and decided to label 14 companies with the largest number of publications as Big Tech.
The whole pipeline of memes extraction is shown in Fig. 10. All used abstracts were preprocessed first – we lowered all words and tokenized them. For the extraction of noun chunks, we used the spaCy library with the en_core_web_md model.
3.3.1 Papers’ affiliation
First, we look at the category of the company as it is specified in the OpenAlex Database. As Big Tech, we take a subset of companies that were among the top 60 companies by market capitalization in the years 2016-2019. The following companies fall into this category: Apple, Google, Microsoft, Facebook, Tencent, Oracle, Intel, IBM, Cisco Systems, TSMC, SAP, Qualcomm, Amazon, Siemens, Alibaba Group, Nvidia, and Samsung. Papers affiliated with Big Tech are 38% of the papers affiliated with the company category, but they account for 57% of their citations.
4 Results
4.1 Non-binary distinction between Big Tech and academy–affiliated papers
We provide substantial arguments, advocating for the less coarse distinction between Academia and Big Tech affiliations than the bisection often used in the literature. The most natural way to follow this proposition would be to use a continuous space by simply calculating for each paper the proportion of Big Tech affiliations to the total number of affiliations. However, as shown in Fig. 2 these values are unevenly distributed – the vast majority are simply academic papers (almost 98% of all papers in the examined dataset), around 0.5% belong only to the Big Tech class, and the remaining part is characterized by a mixed affiliation. The last group also suffers from obvious finite-size effects, i.e., due to a limited number of authors, fractions such as , and tend to occur more frequently than other values (see Fig. 2). To avoid the above obstacles and to be consistent with the observations presented in Fig. 2, we decided to use a ternary affiliation scheme, i.e., Academia–mixed Big Tech–Big Tech.
Let us first explore the implications of the ternary division using simple network metrics: incoming degree and PageRank (Brin and Page, 1998). Both these measures can be used to quantify the importance of a node (paper). In the case of the in-degree, we simply tally the number of citations a paper has received, while PageRank provides more nuanced information. Figure 3 shows box plots of in-degree (panel a) and PageRank (panel d), respectively, grouped according to the introduced ternary affiliation classification. In both cases, the Kruskal-Wallis test (Kruskal and Wallis, 1952) finds significant differences among all groups (p-value ), while post-hoc pairwise comparisons using Dunn’s test (Dunn, 1964) with Holm-Bonferroni corrections (Holm, 1979) indicate that the described distinctions are maintained for all pairs of groups. To prove that the division into three categories does not lead to a loss of additional information available in the continuous approach, we divide the dataset into six categories (Fig. 3c and Fig. 3f): the first and the last are the same as in the previous division, while the remaining four describe cumulative quartiles without exclusive Big Tech affiliations (note that the category labeled as “>0%” in Fig. 3c,f is equivalent to “mixed Big Tech” in Fig. 3a,d). Indeed, also at this level, we obtain similar results for the Kruskal-Wallis test (significant differences with p-value ) as well as for pairwise comparisons, the only exception being the penultimate category (“>75%”), which is indistinguishable from the pure Big Tech affiliations in the case of PageRank.
To examine the reasons for the clear differences between the elements of our ternary classification, we focus on the in-degree and PageRank probability distributions shown in Fig. 3b and Fig. 3e, respectively. Clearly, mixed and Big Tech groups differ from Academia by being characterized by higher values of probability density for larger values of in-degree or PageRank. On the other hand, the difference between mixed and Big Tech is not as pronounced. The likely main reason for the observed differences is revealed when we decide to focus on the number of zero-degree nodes – the proportion of such nodes in Academia, mixed and Big Tech is 0.648, 0.538, and 0.613, respectively. Although the ratio of zero-degree nodes in the mixed category is visibly lower than in the other two categories, we confirm this observation by performing pairwise two-sample proportion tests with Holm-Bonferroni adjustment, which give p-values of for Academia–mixed Big Tech, for Academia – Big Tech, and for mixed Big Tech– Big Tech. An equivalent of Fig. 3 for the network with removed zero-degree (isolated) nodes is shown in D.
Lastly, the decision to introduce the Big Tech category instead of looking at the OpenAlex company category was supported by the intuition that Big Techs differ significantly from other company actors and that extrapolating conclusions from all private companies to Big Tech would be unjustified. To test whether Big Tech are indeed different from the rest of the private companies, we compare the PageRanks of Big Tech–affiliated and company-affiliated papers. We perform the Kruskal-Wallis test between the PageRanks of Big Tech–affiliated papers and papers affiliated with companies that are not Big Tech. The test results confirm that the distribution of PageRanks of Big Tech-affiliated papers is indeed different from that of other companies (p-value < 0.05). Nevertheless, analyzing the ternary division from the company’s point of view (instead of Big Tech) does not bring any additional insights, as can be seen in Fig. 4 (see also D for results obtained with zero-degree nodes removed).
4.2 Most contagious ideas in AI research
Using the meme score measure, we were able to identify and analyze which ideas are the most prevalent in AI research in general, which is the first step in differentiating Big Tech from Academia.
To perform this analysis, we calculated the meme scores, and propagation scores and counted the occurrences for all memes present in our dataset, which resulted in over 1.7 million observations. Such a vast amount of data had to be filtered to get the most interesting results, that is why we decided to limit our study to the phrases where the meme score was greater than 0, which resulted in over 60 000 meaningful observations.
Since our main goal is to deeply analyze the most spreading ideas, we limited the number of memes according to two criteria: the observed occurrences and their meme score values. The first criterion is based on the idea that the memes that have high meme scores but only occur once or twice are not really meaningful topics, as they are too niche to be considered contagious. The second condition is obvious, since the higher the meme score, the more likely it is to spread. To choose a cutoff value, we analyzed the curves shown in Fig. 5. Finally, after limiting ourselves to phrases with a meme score above 0.25 and at least 20 occurrences, we ended up with 251 observations representing the top memes in our dataset.
The set of chosen top memes was later annotated by the group of domain experts which is briefly described in C, resulting in the assignment of one category per observation. The resulting categories are ML algorithms, Medical terms, Security, Niche ML applications, ML Concept, Maths, Graphs, Computer Vision, Natural Language Processing, Data Related, and Other. The annotations of the top memes allow us to analyze which high-level topics are popular among ML researchers, indicating the most promising research areas. According to Table 1, the most popular topic among scientists is the medical applications of ML. This category has the most memes and a third number of the memes occurrences in our dataset. It is not the only application, however, as we can see on the top memes list other important topics like security, and a whole category describing niche applications of ML solutions. However, this doesn’t mean that today’s research focuses only on ML applications, as the second-highest average meme score value, and the second sum of the occurrences belong to the topic describing the ML algorithms. Additionally, we were able to distinguish some particularly popular directions of AI development, such as Computer Vision or NLP. Finally, such an in-depth analysis proves the quality of our framework as the category Other, which includes unrelated topics, as well as meaningless memes, is responsible for only 20% of detected memes. The group has the biggest number of occurrences due to the memes such as ‘paper’, ‘results’, or ‘art’ (part of ‘state of the art’ phrase), which occur multiple times in the majority of the papers.
| Category | Number of memes | Mean meme score | Mean occurrences | Sum of occurrences |
| Medical term | 70 | 0.339 | 141 | 9865 |
| Other | 50 | 0.347 | 2620 | 130993 |
| Niche ML application | 24 | 0.345 | 56 | 1336 |
| Security | 23 | 0.346 | 107 | 2454 |
| ML Algorithm | 20 | 0.380 | 581 | 11616 |
| ML Concept | 20 | 0.343 | 73 | 1456 |
| Computer Vision | 17 | 0.331 | 79 | 1336 |
| Maths | 9 | 0.348 | 35 | 312 |
| Data Related | 7 | 0.405 | 48 | 338 |
| NLP | 7 | 0.332 | 204 | 1430 |
| Graphs | 4 | 0.384 | 47 | 188 |
As we already know the big picture of the main topics in the scientist’s interests, we can combine it with a more detailed analysis, and provide a wider perspective for our top memes group.
Figure 6 shows the characteristics of the top memes subset. As the meme score value depends on the number of memes occurrences as well as the propagation score we can easily imagine that the line that stands for the meme score equal to 0.25 is used as a cutoff value in this analysis. From the density of observations, especially in our group, we can grasp the tendency that the highest meme scores were assigned for the observations with a smaller number of occurrences. Additionally, the plot helps us detect the outliers aforementioned in the previous section, as they yield low propagation score, but also have a large number of occurrences. Thanks to the exemplary meme annotations we can also track some memes and analyze the reasons why are they considered important. For example, ‘fairness’ is considered important mostly due to its high propagation score, whereas ‘cnn’ or ‘deep learning’ benefits from a vast number of occurrences.
To finalize the topic analysis, we decided to present all top memes in the form of a structured word cloud presented in Fig. 7. This way, we are able to analyze particular memes that build each of the predefined categories. We should particularly focus on the medical memes, as this topic is definitely the most important one, due to the abundance of different memes. The memes creating this topic (presented as orange), cover various medical terms. We can discover which disease treatments and studies greatly benefit from the usage of ML (ex. various cancers, epilepsy, diabetic retinopathy, Alzheimer’s disease, malaria, etc.). We can also say that another major points of interest are the studies considering genomics (lncRNA, miRNAs, microRNAs, etc.), and that ML is widely used for various examinations (mammograms, eeg, ecg, wireless capsule endoscopy, etc.).
4.3 Differences in contagiousness between Companies, Academia, and Big Tech
To assess whether contagiousness varies depending on author affiliation, we conducted a comparative analysis of conditioned sticking factors across different affiliation groups. A selection of the 15 highest-ranked memes in each affiliation (Academy, mixed Big Tech, Big Tech, mixed Company, and Company) ordered according to their conditioned sticking factor is shown in Fig. 8a. Although this limited picture suggests that Academia memes dominate, we need to underline that this is simply an illustrative excerpt from the whole distribution.
In order to perform a more systematical analysis, we compared all non-zero conditioned sticking factors across all affiliation groups using Kruskal-Wallis test. Results indicate a significant difference across groups in terms of the sticking factor, i.e. affiliation is a factor that creates distinctions among sticking factor values. Outcomes of post-hoc comparison performed using Dunn’s test between categories presented in Fig. 8b pinpoint distinctions among the sticking factor distributions: strikingly, there are no statistical differences among Academia, Big Tech, and Company. On the other hand, such pairs as mixed Big Tech – Big Tech or mixed Company – Company can be distinguished. This is exactly opposite to the relations observed using network metrics for non-isolated nodes and shown in D, where differences between mixed Big Tech – Big Tech and mixed Company – Company were negligible while one could clearly distinguish Academia, Big Tech, and Company categories.
The above analysis is conducted in absolute terms, i.e., the distributions can be biased by the fact that Academia affiliations are exceedingly more frequent than Big Tech or Company ones (cf Fig. 2). To overcome this issue we focused on memes that appear in at least two different affiliation categories, e.g., “random neural network” (present in mixed Company, Company, mixed Big Tech, and Big Tech) or “captcha” (present in Company, mixed Big Tech, and Big Tech) depicted in Fig. 8a. Following, we conducted Wilcoxon signed-rank tests (Wilcoxon, 1945), comparing sticking factors conditioned on specific groups using matched samples (the same memes in both groups). In case we were able to reject the null hypothesis (observations in both groups and are exchangeable) in favor of the one-sided alternative hypothesis (differences are stochastically larger than a distribution symmetric about zero) we noted it with a directed link in Fig. 8c, pointing from to . The results revealed that memes affiliated with Big Tech exhibited higher sticking factors compared to all other affiliation categories, with company-affiliated memes ranking second. Academic affiliation, on the other hand, was associated with a decrease in the sticking factor compared to other affiliations.
Following this, we explore how this difference depends on the meme category introduced in Table 1. Five hundred memes with the highest meme scores, excluding those not present at least once in papers from each affiliation group, were selected for this comparison. After removing duplicates and the least common categories of memes, the resulting list has 148 memes. The resulting comparison of conditioned sticking factors divided by meme category is presented in Fig. 9.
5 Discussion and conclusions
This work aimed to identify the most contagious ideas (memes) in AI research and assess to what extent the contagiousness of memes depends on the affiliation group of authors of the paper, and whether this contagiousness differs between pure and mixed academic or company affiliation. For this purpose, we analyzed AI papers’ abstracts, affiliations, and citation networks. We employed two key metrics, namely the meme score and the sticking factor, to quantify the contagiousness of these memes. Additionally, we introduced the concept of conditioned sticking factor to assess the relationship between affiliations and contagiousness.
Our findings reveal several noteworthy insights. Firstly, we observed that papers affiliated with Big Tech tend to receive disproportionate citation counts in comparison to all companies. However, the most highly cited papers are those co-authored by individuals from both Big Tech and academic institutions, i.e., papers having mixed affiliations. Additionally, papers co-authored by individuals from both corporate and academic backgrounds exhibit scientometric distinctions from papers authored exclusively by Academia or corporate entities. This pattern holds true for both Big Tech and Company, thus challenging the binary distinctions often assumed in prior research.
The analysis of contagiousness indicates that on average, the contagiousness of Academia-authored memes does not differ from from those authored by companies or Big Tech, although when taking into account only memes present in both affiliations, the same memes have on average higher contagiousness when mentioned by Big Tech than Academia. These facts together could be linked with the difference in the diversity of memes produced and propagated by those two groups. Academia creates far more memes than Big Tech or Companies (mostly due to the difference in overall papers authored), resulting in mean contagiousness being not significantly different than Big Tech or Company. When comparing common ideas of interest of those two groups, the difference in contagiousness becomes apparent.
The impact of affiliation group over the contagiousness of memes could be also understood by applying the framing theory. Framing theory states that an issue can be perceived from various perspectives, each carrying implications for multiple values or considerations (Chong and Druckman, 2007). By applying this theoretical lens to the realm of memetics, we could view issues as memes. In this context, we assert that each meme possesses the potential for diverse framings. These framings, tied to the affiliations of the authors, become integral components that propagate alongside the meme. Consequently, the contagiousness of a meme, when conditioned on a specific group, signifies the extent of that group’s influence over the framing of the respective meme. In essence, the more contagious a meme becomes within a particular group, the more pronounced is that group’s control over the framing associated with the meme. To assess this further, one could analyze the co-occurrence of memes, and ties of such memes “conglomerates” to affiliations.
Further direction of research should also take into account the dynamics of memes contagiousness, investigating whether the impact of affiliation on contagiousness – potentially representing a gap between Academia, Companies, and Big Tech – is expanding or contracting would provide valuable insights.
A limitation of this study, inherent in the original meme score calculating technique, is the operationalization of memes as simple noun chunks, overlooking the potential synonymity of phrases. Consequently, there is a possibility that high-spreading memes may be overlooked. Possible solutions could be techniques such as clustering.
The analysis which we have presented confronts the influence of Big Tech with that of academic institutions on the propagation of ideas between scientific papers. In doing so, we offer a unique perspective on tracing the evolution of ideas influenced by both major technological companies and academic institutions. The conditioned sticking factor, introduced in this work, represents a novel tool that can extend beyond the confines of this study. It can be employed to analyze how various modalities, such as country of origin, journal, or institution, impact the spread of ideas in different contexts.
In conclusion, our results suggest that the prevalent notion of Big Tech’s dominance over AI research is overly simplistic. We advocate for a more nuanced understanding of the roles played by Big Tech companies, corporations, and Academia in shaping the future of AI. We believe that this nuanced perspective, as presented in our work, will contribute to more informed and reasoned discussions about the trajectory of AI research.
5.0.1 Acknowledgements
The research was funded by (POB Cybersecurity and Data Science) of Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. This work was also funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, http://ominoproject.eu). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. Computational part of this study was supported in part by the Poznań Supercomputing and Networking Center (grant number 607).
Appendix A Data gathering and processing
Following Liu et al. [2021] we employed simple search strategy for AI papers. We searched for following words from titles and abstracts: artificial intelligence, machine learning, classifier, neural network, deep learning, data science, nlp, machine-learning, computer vision.
Appendix B OpenAlex categories of affiliations treated as Academia
| Category | Most important institutions |
| Tsinghua University | |
| Shanghai Jiao Tong University | |
| education | Carnegie Mellon University |
| Stanford University | |
| Massachusetts Institute of Technology | |
| Chinese Academy of Sciences | |
| French National Centre for Scientific Research | |
| government | National Institute of Health |
| French Institute for Research in Computer Science and Automation | |
| Commonwealth Scientific and Industrial Research Organisation | |
| Max Planck Society | |
| Electric Power Research Institute | |
| nonprofit | German Research Centre for Artificial Intelligence |
| Instituto de Salud Carlos III | |
| SRI International | |
| Ecole Polytechnique Federale de Lausanne | |
| National Institute of Informatics | |
| facility | Electronics and Telecomunications Research Institute |
| United States Air Force Research Laboratory | |
| Italian Institute of Technology | |
| Mayo Clinic | |
| Boston Children’s Hospital | |
| healthcare | Brigham and Women’s Hospital |
| The University of Texas Southwestern Medical Centre | |
| Vanderbilt University Medical Centre |
Appendix C Annotations
The annotation process was conducted for the top memes, defined as the ones that have a meme score value over 0.25, and more than 20 occurrences, by the team of domain experts. The pipeline for the creation of the annotations consisted of several steps, described below.
1.
Extraction of basic parameters: meme score, meme occurrences, and meme phrases.
2.
Providing a short description of more complex, and not obvious memes. The research was focused on providing the descriptions connected to ML, as we know, that used dataset comes from this domain. Providing the titles of ML papers found connected to researched memes.
3.
Providing the flags for duplicate or very similar memes. Most of them were the abbreviations and their elaborations.
4.
Creation of first frequently repeating topics (ML algorithms, Medical terms, Security, Niche ML application, ML concept, and Other), based on the annotation process.
5.
Iterative extraction of another topic from the Other group (Maths, Graphs, Computer Vision, NLP, Data Related).
6.
Revising whole annotations and re-assignment to proper topics.
Appendix D Network analysis without isolated nodes
Here we perform the same analysis as carried out in Sec. 4.1, however, we restrict the original dataset by removing zero-degree nodes. Regardless of the type of the last category (i.e., Big Tech or company) as well as of the variable in question (in-degree, PageRank) Kruskal-Wallis test finds significant differences among all groups (p-value ). However, as can be seen in Figs 11 and 12 post-hoc tests indicate a lack of difference between mixed category and Big Tech (or company) one, underlining the importance of zero-degree nodes in this context.